OpenStudy (loser66):

How to do row reduce echelon form for this matrix? given \(p^3+q^3+r^3-3pqr \neq 0\) \(A= \left[\begin{matrix}p&q&r|p+q+r\\q&r&p|p+q+r\\r&p&q|p+q+r\end{matrix}\right]\) Please, help.

OpenStudy (loser66):

I tried many ways:). 1) Use manual method 2) find x, y, z by Ax = b --> \(x=A^{-1} b\) to get the solution first, then derive from it to get what term I can manipulate on row reduced echelon method. But it is really a mess. I would like to know the simpler way. Notice: it is a symmetric matrix. and the given equation is -det A

ganeshie8 (ganeshie8):

.

OpenStudy (mathmate):

What did you get by manual method?

OpenStudy (loser66):

I got nothing!! That is why I have to derive from method 2

OpenStudy (loser66):

Let T = p+q+r and we have A is the 3x3 matrix from the LHS \(det A = -p^3-q^3-r^3 +3pqr = (-1) (p^3+q^3+r^3-3pqr)\neq 0\) Hence, A has inverse. By using cofactor transpose of matrix A, I got \(A^{-1}=\dfrac{-1}{det A}\left[\begin{matrix}rq -p^2&pr-q^2&pq-r^2\\rp-q^2&pq-r^2&qr-r^2\\qp-r^2&qr-p^2&pr-q^2\end{matrix}\right]\)

OpenStudy (loser66):

Hence \[\vec x=A^{-1}b =\dfrac{-T}{det A}\left[\begin{matrix}rq -p^2&pr-q^2&pq-r^2\\rp-q^2&pq-r^2&qr-r^2\\qp-r^2&qr-p^2&pr-q^2\end{matrix}\right]\]

OpenStudy (loser66):

After simplifying, I got \(x_1= 1+\dfrac{qr^2+q^2r-pr^2-p^2r}{p^3+q^3+r^3-3pqr}\)

OpenStudy (loser66):

From here, I argue that, to get rref , we need entry \(a_{11} \) must be the form of \(\dfrac{p^3+q^3+r^3-3pqr}{p^3+q^3-r^3-3pqr}=1\) and the rhs, as above, So that we can have the required solution. However, I have no way to get \(a_{11}\) as above. Or, I made mistake at somewhere. That is what I tried.

ganeshie8 (ganeshie8):

\(A= \left[\begin{matrix}p&q&r|p+q+r\\q&r&p|p+q+r\\r&p&q|p+q+r\end{matrix}\right]\)

OpenStudy (mathmate):

Indeed it's messy. If you find the row echelon for of A (3x4), you will find that the last row is [0,0,1,1], which means that r=1. By symmetry of the three variables, we can conclude that p=q=r. However, by the same token, as long as p=q=r=k, k can take on any real value. So we have an infinite number of solutions!

OpenStudy (mathmate):

Oops, did not read the first line. This means that symmetric solutions are out.

OpenStudy (loser66):

nope, if p=q=r =1, then \(p^3+q^3+r^3-3pqr=0\) contradict

OpenStudy (mathmate):

The row echelon form turns out to be [1, q/p, r/p,(p+q+r)/p], [0, 1, (p^2-qr)/(pr-q^2), ((p-q)r+p^2-q^2)/(pr-q^2)], [0, 0, 1, 1] But I have no idea where to go from here. @ganeshie8

OpenStudy (loser66):

We cannot divide by any p, q, r since we don't know whether they are =0 or not.

OpenStudy (mathmate):

@Loser66 To investigate further, we have to invert a=matrix A less the augmented part =[p,q,r] [q,r,p] [r,p,q] and the inverse of \(\large a^{-1}=\left[\begin{matrix}\frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\\ \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\\ \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\end{matrix}\right]\) and \(a^{-1}.\left[\begin{matrix}p+q+r\\p+q+r\\p+q+r\end{matrix}\right]\) =\(\left[\begin{matrix}1\\1\\1\end{matrix}\right]\) Which means that any combination of p,q,r such that p+q+r=1, and \(p^3+q^3+r^3-3pqr \neq 0\) would fit the bill. For example, (1/4,1/4,1/2), or even (-1/4,-1/4,1) are solutions.

OpenStudy (loser66):

I got (1,1,1) also for x, y , z, not for p, q, r the third way I approach is \(p^3+q^3+r^3-3pqr=(p+q+r)*(p^2+q^2+r^2-pr-pq-rq)\) the LHS\(\neq 0\), hence \((p+q+r)\neq 0 ~~and~~~ (p^2+q^2+r^2)\neq 0\)

OpenStudy (loser66):

So that to A, to get rref, I do \((R_1+R_2+R_3) \) for all Rows that gives me all Rows are the same and each entries for the rows is \((p+q+r) ~~~(p+q+r) ~~~ (p+q+r) | 3 (p+q+r)\)

OpenStudy (loser66):

Then multiple that row by \((p^2+q^2+r^2-pr-pq-rq)\) and simplify we have all entries are \(p^3+q^3+r^3-3pqr \), the right hand side is \(3 (p^3+q^3+r^3-3pqr)\)

OpenStudy (loser66):

divide the rows by its common entry, we have (1 1 1 = 3) And we get 1x + 1y+ 1z = 3 That is x=3-y-z

OpenStudy (mathmate):

@loser66 What's the difference between x,y,z and p,q,r. Do they solve the same problem? Sorry, I don't get what you mean by: "I got (1,1,1) also for x, y , z, not for p, q, r"

OpenStudy (loser66):

The original problem : Solve the following system of equation for x, y, and z in \(\mathbb R^3\), by row reduction and check by substitution that your solution is correct. px + qy + rz = p+q+r qx + ry+pz = p+q+r rx + py + qz = p+q+r It is given that \(p^3+q^3+r^3-3pqr \neq 0\) Do not divide by p, q, or r, as you do not know if any of these equals 0 Write the answer in 1) vector-parametric form 2) flat form 3) functional form 4) What happens to the solution set if \(p^3+q^3+r^3-3pqr=0\)

OpenStudy (mathmate):

@loser66 Thank you for posting the complete original question. That's a little more clear to me now. I was confused earlier, because the x,y,z were understood in the matrix A, and it did not occur to me. I have posted the following in an earlier reply, but not in fancy formatting as the following, so it's worthwhile to repeat as follows: the row echelon form of A is: \[\left[\begin{matrix}1 & \frac{q}{p} & \frac{r}{p} & \frac{p+q+r}{p}\\ 0 & 1 & \frac{p^2-qr}{pr-q^2} & \frac{(p-q)r+p^2-q^2}{pr-q^2}\\ 0 & 0 & 1 & 1\end{matrix}\right]\] which by normal backward pass gives (x,y,z)=(1,1,1). As a check, let a= square matrix formed by the first three columns of A, i.e. a=\(\left[\begin{matrix}p & q & r\\ q & r & p\\ r & p & q\end{matrix}\right]\) the inverse of a is \( a^{-1}=\left[\begin{matrix}\frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\\ \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\\ \frac{pq-r^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{qr-p^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)} & \frac{pr-q^2}{r(pq-r^2)+p(qr-p^2)+q(pr-q^2)}\end{matrix}\right]\) and \(\left[\begin{matrix}x\\y\\z\end{matrix}\right]=a^{-1}.\left[\begin{matrix}p+q+r\\p+q+r\\p+q+r\end{matrix}\right]=\left[\begin{matrix}1\\1\\1\end{matrix}\right]\) exactly the same solution as by row echelon form. You may notice that the denominator of the inverse is exactly \(p^3+q^3+r^3-3pqr\), which therefore calls for the restriction given in the original question.